OpenStudy (anonymous):

whats the simple formula for the nth term. form an+b. 10, 14, 18,22, 26....

OpenStudy (anonymous):

a1=10 a2=14 a2-a1=d 14-10=d d= 4 an+1= an +d

OpenStudy (anonymous):

Each term is increasing by 4. So, nth term is 4n + 6.

OpenStudy (cwrw238):

nth term = 10 +4 (n-1)

OpenStudy (anonymous):

It said to put it in the form of "an + d". So, it is 4n + 6.

OpenStudy (cwrw238):

which simplifies to what GT said